Could you help me solve this problem?
if y varies inversely as x^2, and y=2 when x=4, what is the constant of the variation? 2?

Thank you.
: Could you help me solve this problem?
if y varies inversely as x^2, and y=2 when x=4, what is the constant of the variation? 2?

Thank you.

Answer by ankor@dixie-net.com(4667)

If y varies inversely as x^2," equation for this is:
yk/x^2; inversely means, when x increases, y decrease
:
y=2 when x=4, what is the constant of the variation?
:
Substituting for y and x we have:
2 = k/4^2
2 = k/16
Multiply both sides by 16
2*16 = k
:
k = 32; (the constant of variation)
:
:
Check this: using k=32 and x=4
y = 32/x^2
y = 32/16
y = 2